If the density of methanol is 0.793 kg L–1, what is its volume needed for making 2.5 L of its 0.25 M solution?

Grade:12

3 Answers

Gaurav
9 years ago

Molar mass of methanol (CH3OH) = (1 × 12) + (4 × 1) + (1 × 16)= 32 g mol–1= 0.032 kg mol–1 (divide by 1000 toconvertin Kg )
Molarity of methanolsolution=0.793 Kg L-1/ 0.032 Kg mol -1= 24.78 mol L–1

(Since density is mass per unitvolume)
Applying,
M1V1 = M2V2
(Givensolution) (Solutiontobe prepared)
(24.78 mol L–1) V1= (2.5 L) (0.25 mol L–1)V1
V1 = 0.0252 L (multiply by 1000 toconvertin ml )
= 25.22 mL
Suresh
5 years ago
25.22 ml vvggbfxshkhhgtgogyydgtijgrt FedEx grdcffd drdfgvb tghjtfb ttghyhbvv fgutdyh ggfsdy dfgghub     fhfhgjyvn  fvfhjehuov   yjhgjuhsfhh    fhjhjgyjiv   ghhjoojj     2446 677678999787764566 45767877888887766889998   785678877888875578
Swati
5 years ago
Molar mass of methanol (CH3OH)=12+3*1+16+1=32g mol-1
32*1000 =0.032kgmol-1      (to convert gram into kilogram)
Molarity of methanol=0.793kg/0.032kgmole-1=24.78mole-1  (density =mass/volume)
M1 V1 =M2V2
24.8 V1 =2.5*0.25
V1=0.0252L
V1=25.22ml . (Convert l to ml multiply 1000)